An investment of R20 000 accumulated to R45 200. If the applicable simple interest rate is 12% per year, then the time under consideration is

1.     10,50 years.

2.     15,12 years.

3.     7,19 years.

4.     4,65 years.

5.     3,25 years.


1
Expert's answer
2021-05-14T14:46:07-0400

"Interest=principal \u00d7rate \u00d7time"

"Interest=amount-principal"

Interest"=45200-20000=25200"

"25200=20000×\frac{12}{100}×t"

"25200=2400t"

"t=10.50\space years"

Hence option 1 is correct.